drug response curve describes the level of medication in the bloodstream after a drug is administered. A surge function S(t) = Atpekt

 (where t > 0) is often used to model the response curve, reflecting an initial surge in the drug level and then a more gradual decline. If, for a particular drug, A = 0.03, p = 4, k = 0.05, and t is measured in minutes, estimate the times t corresponding to the inflection points. (Round your answers to two decimal places.)t = min (smaller value) t = min (larger value)


1
Expert's answer
2022-03-24T13:30:08-0400

"S(t)=At^pe^{-kt}=0.03\cdot t^4\cdot e^{-0.05t}"; "t>0"

Find inflection points.

Solution:

Inflection points can only occur when the second derivative is zero or undefined.

Here we have:

"S'(t)=0.03\cdot 4t^3e^{-0.05t}-0.03\cdot 0.05t^4e^{-0.05t}=""0.03e^{-0.05t}t^3( 4-0.05t)";

"S''(t)=-0.03\cdot 0.05e^{-0.05t}t^3( 4-0.05t)+" "0.03\cdot 3e^{-0.05t}t^2( 4-0.05t)-""0.03\cdot 0.05e^{-0.05t}t^3=" "t^2\cdot e^{-0.05t}( 0.000075t^2-0.012t+0.36)";

"t^2\cdot e^{-0.05t}( 0.000075t^2-0.012t+0.36)=0"

Therefore possible inflection points occur at

"t=40" and "t=120"; ("t=0" not interested because we are looking for inflection points only for "t>0" ).

However, to have an inflection point we must check that the sign of the second derivative is different on each side of the point. Here we have

"S''(10)\approx 15.01";

"S''(100)\approx-6.06";

"S''(150)\approx3.08".

Hence, both are inflection points.

Answer: there are two inflection points "t=40".00 and "t=120".00 .